# Autotune: fast, accurate, and automatic tuning parameter selection for LASSO

### v1

@@ 2 Background
Recently [kock2025data] proposed an automated tuning mechanism for VAR Lasso, building upon the iterative algorithm proposed in [belloni2012sparse]. Denote $\widehat{\Phi}^{(t)}$ as the estimate of $\Phi$ at the $t^{th}$ iteration. Then, [kock2025data] iteratively computes the following columnwise weighted lasso estimator for $t=1,2,3,\ldots$
$\widehat{\Phi}^{(t)}_{\cdot,i}=\underset{\beta\in\mathbb{R}^{dp}}{\operatorname{argmin}}\frac{1}{n}\mathopen{}\mathclose{{\left\|\text{Y}_{\cdot,i}-\text{X}\beta}}\right\|^{2}+\frac{\lambda}{n}\mathopen{}\mathclose{{\left\|{\Gamma^{(t)}_{i}\beta}}}\right\|_{1},\quad\text{ for }i=1,\ldots,p$ (23)
@@
Coordinate descent’s ease of implementation, availability of coordinate-wise closed-form updates, augmented with all the computational tricks of warm starts, and active set selection makes it the default and the fastest algorithm [10.1214/07-AOAS131] for researchers using Lasso over other algorithms like proximal gradient descent [nesterov2013introductory], and LARS [efron2004least]. Additionally, it enjoys convergence guarantees [tseng2001convergence]. CD’s idea is first to fix the $\lambda$ in ˜2 and then iteratively optimize over each parameter holding other parameters constant at their current values.
@@ 1 Initialize $\widehat{\beta}^{(0)}=0_{p}$, $r=$ $\text{X}\widehat{\beta}^{(0)}=Y$
We provide the pseudocode of Lasso at a fixed $\lambda$ in Algorithm 1. A core component of CD updates in Lasso is the partial residual $r^{(j)}$, $\forall j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$, which comes out to be $r^{(j)}=r+\text{X}_{.,j}^{\top}\widehat{\beta}_{j}$ (from ˜4), where $r=\text{Y}-\sum_{k\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}}\text{X}_{.,k}^{\top}\widehat{\beta}_{k}$ is the full residual. Partial residuals are continuously computed and updated inside every CD step, but there has been only a limited discussion about it in the prior literature. We take a deeper look into the partial residuals in the next section.
@@
Sometimes, to further reduce computational costs, a subset of coefficients called “active set” is selected via some tests [10.1111/j.1467-9868.2011.01004.x, hastie2015statistical] at each $\lambda$ in the grid to gather all the predictors whose absolute correlation with the YY/residuals is not below a certain threshold depending on $\lambda$. This ensures that the predictors will contribute to the prediction of $Y$ at that $\lambda$. Coefficients outside the active set have the value 0, and the CD algorithm is iterated only over the active set.
In the state-of-the-art algorithms, scaled Lasso internally uses LARS, which is computationally slower than CD [10.1214/07-AOAS131], and K-fold CV Lasso computes the entire Lasso solution path $K+1$ times. autotune Lasso runs the coordinate descent(CD) once at each $\lambda$ in its smaller regularization path(than CV lasso, see Fig.˜4). Therefore, we may therefore expect $\mathsf{autotune}$ to be faster (Sec.˜4.1.3, 4.2).
Sequential F-tests.
Sequential F-tests/sum of squares [Chap 4.6, montgomery2020introduction] is a classic model selection tool which is often employed to calculate ANOVA tables [carpackage]. Given the responses Y and $p$ predictors Xj, $j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$, it looks into the following sequence of nested linear models and their corresponding residual sum of squares (RSS):
$\text{RSS}(\mathcal{M}_{k})$ $=\sum_{i=1}^{n}\mathopen{}\mathclose{{\left(\text{y}_{i}-\widehat{\text{y}}_{i}}}\right)^{2}\quad\text{where $\widehat{\text{y}}_{i}$'s comes from fitting $\mathcal{M}_{k}$ to data.}\quad$ (26)
@@
until either we arrive at a $k_{0}\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$ where we fail to reject the $\mathcal{H}_{0,k_{0}}$ or we reject $\mathcal{H}_{0,j}$ $\forall j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$. If the procedure stops, we conclude that the significant predictors for the regression of Y on X are $\mathopen{}\mathclose{{\left\{1,\dots,k_{0}-1}}\right\}$, otherwise none of the predictors are spurious. Note that in $n<p$ regime, we will always arrive at a $k_{0}\leq n$ where the sequential F-test stops. The rank order of predictors in which the nested models $\mathcal{M}_{k}$ are constructed is important for selecting the final model.
@@ 3 Method: tuning with $\mathsf{autotune}$
In our empirical analyses, we found that these strategies of iteratively updating $\sigma$ suffer from the shrinkage bias in $\widehat{\beta}$. So we develop a more aggressive bias correction strategy based on the partial residuals (PR) which are already being computed inside Algorithm 1.
A basic autotune strategy.
@@
Note that $\epsilon^{(k)}$ is essentially the vector of residuals from the linear model $\mathcal{M}_{k-1}$ of $Y$ on $\mathopen{}\mathclose{{\left\{\text{X}_{\cdot,1},\text{X}_{\cdot,2},\dots}}\right\}$. Then, we can perform F-test of $\mathcal{M}_{k-1}$ against $\mathcal{M}_{k}$ (the linear model we get after adding $u_{\cdot,k}$ to $\mathcal{M}_{k-1}$) as follows:
$\mathrm{RSS}(\mathcal{M}_{k})$ $=\mathopen{}\mathclose{{\left\|{\epsilon^{(k)}}}}\right\|_{2}^{2}\quad\text{and}\quad\mathrm{RSS}(\mathcal{M}_{k-1})=\mathopen{}\mathclose{{\left\|{\epsilon^{(k-1)}}}}\right\|_{2}^{2}$ (30)
and we use the formula in ˜27. Let $k_{0}-1$ be the number of predictors in the final linear model. This procedure reduces the computational complexity to $\mathcal{O}(n.k_{0})$ from $\mathcal{O}(n.k_{0}^{2})$ in the basic procedure where each model $\mathcal{M}_{k}$ is fully fitted without using the previous fit $\mathcal{M}_{k-1}$. The full details of this $\sigma$ update is given in Algorithm 2. So, this version of $\mathsf{autotune}$ is effectively using Algorithms 1 and 2 alternately to perform steps ˜(M1) and ˜(M2) above.
Early stopping in Algorithm 1.
@@
16 [2pt] if $\mathrm{F-test}$ $\mathrm{statistic}\leq\mathrm{cutoff}$ then
@@
19 Include $j^{th}$ predictor $\text{X}_{\cdot,j}$ in the $\mathsf{Support.Set}$
@@
Output: new noise variance estimate $\widehat{\sigma}^{2}$, updated support set $\mathsf{Support.Set}$, $\mathsf{Predictor.Ranking}$.
Algorithm 2 Sigma Update – Updates the estimate of noise variance $\widehat{\sigma}^{2}$.
@@ 5  Set $\lambda=\lambda_{0}\widehat{\sigma}^{2}$
Output: regression coefficients’ estimate $\widehat{\beta}$, noise variance estimate $\widehat{\sigma}^{2}$, regularization parameter $\lambda$
Algorithm 3 Autotune Lasso – Return estimates of regression coefficients $\beta$.
Warm starts and initialization of $\lambda_{0}$.
We use a warm start strategy as is common in glmnet. This means we initialize the algorithm for $\lambda_{k+1}$ at the solution obtained at the previous candidate value of $\lambda_{k}$, for every $k=1,2,\ldots$. Even though our algorithm searches over a sequence of $\lambda$ values which are not necessarily close, we find that this strategy helps with quick convergence of CD as opposed to fixed initializers such as 0.
@@
Prior literature [glmnet, hastie2015statistical] recommends initiating computation of Lasso path at $\lambda=\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}$ which is just big enough so that $\widehat{\beta}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}}}\right)=0$ (recall $\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda}}\right)=\mathop{\mathrm{argmin}}_{\beta\in\real^{p}}\mathcal{L}_{\lambda}(\beta)$). But in $\mathsf{autotune}$, we found out this approach makes the regression estimates $\widehat{\beta}$ stuck at 0. We started seeing the empirical success of $\mathsf{autotune}$ after beginning the Lasso path at a $\lambda$ slightly smaller than $\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}$. Moreover, as $\mathsf{autotune}$ starts from $\widehat{\beta}=0$, we set $\widehat{\sigma}^{2}_{\mathsf{(initial)}}=$Var(Y). Consequently, our final choice of $\lambda_{0}$ ensures that the initial $\lambda=\lambda_{0}\widehat{\sigma}^{2}_{\mathsf{(initial)}}=\frac{1}{2}\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}$.
Speeding up predictor ranking.
@@
Switching from standard deviation, which is essentially the $\ell_{2}$ norm of PR, to an $\ell_{1}$-norm of PR helps improve speed without comprising accuracy. In other words, important predictors are still ranked high in terms of both norms.
Active set selection.
@@
We can also utilize active set selection [10.1111/j.1467-9868.2011.01004.x] to reduce the computational load further. We know from the Lasso’s KKT conditions [hastie2015statistical] that for all predictors $\text{X}_{\cdot,j}$ with nonzero estimates of coefficients, $\mathopen{}\mathclose{{\left|\langle{\text{X}_{\cdot,j}},{\text{Y}-\text{X}\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda}}\right)}\rangle/n}}\right|=\lambda$ and the remaining predictors will have smaller absolute correlation. A reasonable expectation is that at the start of CD at a $\lambda$, predictors with a small inner product with the residuals are more likely to have a zero coefficient as compared to the predictors with bigger inner products. We use this intuition to remove the spurious predictors, which can potentially boost the speed of the computations while still arriving at the optimal minimizer (details in Active-Autotune-Lasso, and its empirical performance is demonstrated in Sec.˜A.2).
3.1 Autotune for VAR Lasso
@@
Estimating a VAR($d$) model on a $p$-dimensional time series under $\ell_{1}$ penalty is equivalent to solving $p$ separate lasso regression problems (Sec.˜2, ˜22). For a $p$-dimensional stationary time series $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}=\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left(Z^{t}_{1},\dots,Z^{t}_{p}}}\right)}}\right\}_{t\geq 1}$, eqn. ˜21 shows how we construct the response and predictor matrix Y and X from $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}$ for estimating $\Phi$, the matrix formed by stacking $d$ VAR transistion matrices. $\mathsf{Autotune}$ VAR(d) performs $p$ separate $\mathsf{autotune}$ Lasso regression of each column of Y on X to yield $\widehat{\Phi}$.
@@
3.2 Noise estimation and sparsity diagnostics with $\mathsf{autotune}$
The intermediate PR and $\widehat{\sigma}^{2}$ of our algorithm show promising results for downstream tasks in high-dimensional regression. In Sec.˜4.3, we show two such examples. First, our $\widehat{\sigma}^{2}$ shows competing accuracy as the state-of-the-art alternatives, and seems to help with high-dimensional inference with bias corrected Lasso. Second, the rank list of PR standard deviations seems to be a useful to diagnose if the underlying model is sparse. We track the improvement in cumulative and adjusted R2 as we build our nested sequence of linear models, and plot these to see any obvious elbow pattern as in scree plot for principal component analysis.
3.3 An illustration of PR properties
Figure 3: Partial residuals’ SD of first 2 coordinate descents (CDs) of $\mathsf{autotune}$ Lasso, which requires 30 iterations of CDs to converge for a high-dimensional setup where $n=80,p=750$. We denote significant predictors (predictors whose true coefficients are non-zero) with “Sig Beta” labeling and green bars, and insignificant predictors are denoted by blue and maroon bars based on whether their coefficients’ estimates are zero/nonzero at the end of that iteration. We also cap the number of insignificant predictors in blue group (their coefficients estimated as zero) to 10 for readability of the plot. Green and blue dashed vertical lines indicates SD of the true residuals and noise SD estimates at the end of the iteration.

### v2

@@ 2 Background
Recently [kock2025data] proposed an automated tuning mechanism for VAR Lasso, building upon the iterative algorithm proposed in [belloni2012sparse]. Denote $\widehat{\Phi}^{(t)}$ as the estimator of $\Phi$ at the $t^{th}$ iteration. Then, [kock2025data] iteratively computes the following columnwise weighted Lasso estimator for $t\geq 1$
$\widehat{\Phi}^{(t)}_{\cdot,i}=\underset{\beta\in\mathbb{R}^{dp}}{\operatorname{argmin}}\hskip 2.84544pt\frac{1}{n}\mathopen{}\mathclose{{\left\|\text{Y}_{\cdot,i}-\text{X}\beta}}\right\|^{2}+\frac{\lambda}{n}\mathopen{}\mathclose{{\left\|{\Gamma^{(t)}_{i}\beta}}}\right\|_{1},\quad\text{ for }i=1,\ldots,p$ (23)
@@
The ease of implementing coordinate descent, availability of coordinate-wise closed-form updates, augmented with all the computational tricks of warm starts, and active set selection makes it the default and the fastest algorithm [10.1214/07-AOAS131] for researchers using Lasso over other algorithms like proximal gradient descent [nesterov2013introductory] and LARS [efron2004least]. The idea of CD is first to fix the $\lambda$ in ˜2 and then iteratively optimize over each parameter holding other parameters constant at their current values.
@@ 1 Initialize $\widehat{\beta}^{(0)}=0_{p}$, $r=$ $\text{X}\widehat{\beta}^{(0)}=Y$
We provide the pseudocode of Lasso at a fixed $\lambda$ in Algorithm 1. A core component of CD updates in Lasso is the PR $r^{(j)}$, $\forall j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$, which comes out to be $r^{(j)}=r+\text{X}_{.,j}^{\top}\widehat{\beta}_{j}$ (from ˜4), where $r=\text{Y}-\sum_{k\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}}\text{X}_{.,k}^{\top}\widehat{\beta}_{k}$ is the full residual. PR are continuously computed and updated inside every CD step, but there has been only a limited discussion about it in the prior literature. We take a deeper look into the partial residuals in the next section.
In Lasso, the standard practice [hastie2015statistical] is to calculate $\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda}}\right)$ over a decreasing $\lambda$ grid of equally spaced values (on log scale) from $\lambda_{\max}$ to (almost) 0. $\lambda_{\max}=\max_{j\in[p]}\frac{1}{n}\langle{\text{Y}},{\text{X}_{.,j}}\rangle$ is the smallest $\lambda$ which ensures $\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda}}\right)=0$. The Lasso solution for each successive $\lambda$ in the grid is then obtained by initializing the coordinate descent (CD) algorithm with the solution from the previous $\lambda$, a technique known as warm starts [hastie2015statistical]. This process is repeated until the smallest $\lambda$ in the grid is reached, yielding $\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda_{\min}}}\right)$. This entire procedure is often called pathwise coordinate descent, and the set of solutions $\mathopen{}\mathclose{{\left\{\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda_{\max}}}\right),\dots,\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda_{\min}}}\right)}}\right\}$ is interchangeably called solution path or regularization path. Warm starts accelerates the solving of pathwise coordinate descent as compared to running CD separately across different values of $\lambda$ in the grid.
Sometimes, to further reduce computational costs, a subset of coefficients called “active set” is selected [10.1111/j.1467-9868.2011.01004.x, hastie2015statistical] at each $\lambda$ in the grid. In contrast to pathwise CD, Scaled Lasso internally uses least angle regression (LARS), which is slower than CD [10.1214/07-AOAS131], and K-fold CV Lasso computes the entire Lasso solution path K+1+1 times. So, a data-driven choice of $\lambda$ values, coupled with running CD in each $\lambda$, has the potential to reduce runtime over CV Lasso and Scaled Lasso.
Sequential F-tests:
Sequential F-tests [Chap 4.6, montgomery2020introduction] is a classic model selection tool which is often employed to calculate ANOVA tables [carpackage]. Given the response Y and $p$ predictors Xj, $j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$, it looks into the following sequence of nested linear models and their corresponding residual sum of squares (RSS):
@@
until either we arrive at a $k_{0}\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$ where we fail to reject the $\mathcal{H}_{0,k_{0}}$ or we reject $\mathcal{H}_{0,j}$ for all $j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$. If the procedure stops, we conclude that the significant predictors for the regression of Y on X are $\mathopen{}\mathclose{{\left\{1,\dots,k_{0}-1}}\right\}$, otherwise all the predictors are included in the model. Note that in $n<p$ regime, we will always arrive at a $k_{0}\leq n$ where the sequential F-test stops. The rank order of predictors in which the nested models $\mathcal{M}_{k}$ are constructed is important for selecting the final model.
@@ 3 Method: tuning with $\mathsf{autotune}$
In our empirical analyses, we found that these strategies of iteratively updating $\sigma$ suffer from the shrinkage bias in $\widehat{\beta}$. So we develop a more aggressive bias correction strategy based on the PR which are already being computed inside Algorithm 1.
A basic $\mathsf{autotune}$ strategy:
@@
Note that $\epsilon^{(k)}$ is essentially the vector of residuals from the linear model $\mathcal{M}_{k-1}$ of $Y$ on $\mathopen{}\mathclose{{\left\{\text{X}_{\cdot,1},\text{X}_{\cdot,2},\dots}}\right\}$. Then, we can perform F-test of $\mathcal{M}_{k-1}$ against $\mathcal{M}_{k}$ (the linear model we get after adding $u_{\cdot,k}$ to $\mathcal{M}_{k-1}$) by using the formula in ˜27 with $\mathrm{RSS}(\mathcal{M}_{k})=\mathopen{}\mathclose{{\left\|{\epsilon^{(k)}}}}\right\|_{2}^{2}$ and $\mathrm{RSS}(\mathcal{M}_{k-1})=\mathopen{}\mathclose{{\left\|{\epsilon^{(k-1)}}}}\right\|_{2}^{2}$. Let $k_{0}-1$ be the number of predictors in the final linear model. This procedure reduces the computational complexity to $\mathcal{O}(nk_{0})$ from $\mathcal{O}(nk_{0}^{2})$ in the basic procedure where each model $\mathcal{M}_{k}$ is fully fitted without using the previous fit $\mathcal{M}_{k-1}$. The full details of this $\sigma$ update is given in Algorithm 2. So, this basic version of $\mathsf{autotune}$ is effectively using Algorithms 1 and 2 alternately to perform steps ˜(M1) and ˜(M2) above.
Early stopping in Algorithm 1:
@@
16 [2pt] if $\mathrm{F}$-$\mathrm{test}$ $\mathrm{statistic}\leq\mathrm{cutoff}$ then
@@
19 $\mathsf{Support.Set}\leftarrow\mathsf{Support.Set}\cup\{j\}$
@@
Output: $\widehat{\sigma}^{2}$, updated support set $\mathsf{Support.Set}$, $\mathsf{Predictor.Ranking}$
Algorithm 2 Sigma Update – Updates the estimate of noise variance $\widehat{\sigma}^{2}$
@@ 5  Set $\lambda=\lambda_{0}\widehat{\sigma}^{2}$
Output: $\widehat{\beta}$, $\widehat{\sigma}^{2}$, $\lambda$
Algorithm 3 Autotune Lasso – Return estimates of regression coefficients $\beta$
Warm starts and initialization of $\lambda_{0}$:
We use a warm start strategy as is common in glmnet. This means we initialize the algorithm for $\lambda_{k+1}$ at the solution obtained at the previous candidate value of $\lambda_{k}$, for every $k\geq 1$. Even though our algorithm searches over a sequence of $\lambda$ values which are not necessarily close, we find that this strategy helps with quick convergence of CD as opposed to fixed initializers such as 0.
@@
Prior literature [glmnet, hastie2015statistical] recommends initiating computation of Lasso path at $\lambda=\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}$ which is just big enough so that $\widehat{\beta}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}}}\right)=0$. But in $\mathsf{autotune}$, we found out this approach makes the regression estimates $\widehat{\beta}$ stuck at 0. We started seeing the empirical success of $\mathsf{autotune}$ after beginning the Lasso path at a $\lambda$ slightly smaller than $\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}$. Moreover, as $\mathsf{autotune}$ starts from $\widehat{\beta}=0$, we set $\widehat{\sigma}^{2}_{\mathsf{(initial)}}=$Var(Y). Consequently, our final choice of $\lambda_{0}$ ensures that the initial $\lambda=\lambda_{0}\widehat{\sigma}^{2}_{\mathsf{(initial)}}=\frac{1}{2}\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}$.
Speeding up predictor ranking:
@@
Switching from standard deviation, which is essentially the $\ell_{2}$ norm of PR, to an $\ell_{1}$ norm of PR helps improve speed without comprising accuracy. In other words, important predictors are still ranked high in terms of both norms.
Active set selection:
@@
We can also utilize active set selection [10.1111/j.1467-9868.2011.01004.x] to reduce the computational load further. We know from the KKT conditions of Lasso [hastie2015statistical] that for all predictors $\text{X}_{\cdot,j}$ with nonzero estimates of coefficients, $\mathopen{}\mathclose{{\left|\langle{\text{X}_{\cdot,j}},{\text{Y}-\text{X}\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda}}\right)}\rangle/n}}\right|=\lambda$ and for the remaining predictors, the previous term will have a value less than $\lambda$. A reasonable expectation is that at the start of CD at a $\lambda$, predictors with a small inner product with the residuals are more likely to have a zero coefficient as compared to the predictors with bigger inner products. We use this intuition to remove the spurious predictors, which can potentially boost the speed of the computations while still arriving at the optimal minimizer (details in Active-Autotune-Lasso, and its empirical performance is demonstrated in Sec.˜A.2). We find that for correlated predictors, this strategy is not very accurate, so in this work we report results without active set selection and leave this for future work.
3.1 Using $\mathsf{autotune}$ for VAR Lasso:
@@
Estimating a VAR($d$) model on a $p$-dimensional time series under $\ell_{1}$ penalty is equivalent to solving $p$ separate Lasso regression problems (Sec.˜2, ˜22). For a $p$-dimensional stationary time series $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}=\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left(Z^{t}_{1},\dots,Z^{t}_{p}}}\right)}}\right\}_{t\geq 1}$, eqn. ˜21 shows how we construct the response and predictor matrix Y and X from $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}$ for estimating $\Phi$, the matrix formed by stacking $d$ VAR transistion matrices. VAR Lasso with $\mathsf{autotune}$ performs $p$ separate $\mathsf{autotune}$ Lasso regression of each column of Y on X to yield $\widehat{\Phi}$.
@@
3.2 Noise estimation and sparsity diagnostics with $\mathsf{autotune}$:
The intermediate PR and $\widehat{\sigma}^{2}$ of our algorithm show promising results for downstream tasks in high-dimensional regression. In Sec.˜4.3, we show two such examples. First, our $\widehat{\sigma}^{2}$ shows competing accuracy as the state-of-the-art alternatives, and seems to help with high-dimensional inference with bias corrected Lasso. Second, the rank list of PR standard deviations seems to be a useful to diagnose if the underlying model is sparse. We track the improvement in cumulative and adjusted R2 as we build our nested sequence of linear models, and plot these to check for any obvious elbow pattern as in scree plot for principal component analysis.
3.3 An illustration of PR properties:
Figure 3: SD of PR of first 2 CD of $\mathsf{autotune}$ Lasso, which requires 30 iterations of CDs to converge for a high-dimensional setup ($n=80,p=750$). We denote significant predictors (predictors whose true coefficients are non-zero) with “Sig Beta” labeling and green bars, and other predictors are denoted by blue and maroon bars based on whether their coefficients’ estimates are zero or nonzero at the end of that iteration. We also cap the number of insignificant predictors in blue group to 10 for readability of the plot. Green and blue dashed vertical lines indicate SD of the true residuals and noise SD estimates at the end of the iteration.